Hawai'i Earthquake Update for Gemini North
(as of 5:00 PM HST, Thursday, November 9, 2006)
A fairly serious earthquake happened Sunday morning near Waikoloa, on the west side of the Big Island. At magnitude 6.6 this was a much stronger earthquake than we are accustomed to here in Hawai'i. Shortly after the event Gemini had staff on the summit to assess the damage to the telescope. We will provide updates as they warrant and information becomes available.
The following item was updated at 5h00 pm HST on Thursday, November 9th
Repairs to the Gemini North Secondary System have been successfully completed! The system was thoroughly tested in the lab and will be reattached to the top end of the telescope on Friday November 10th. The first half of Friday night is currently scheduled for engineering tests and, if these are successful, the telescope will be handed over for instrument tests.
In light of the fact that other Mauna Kea telescopes have run into technical "glitches" once back on the sky we will not be surprised if we find similar small problems once systems are fully exercised.
The following item was updated at 5h00 pm HST on Thursday, October 26
Message from Dr. Doug Simons, Gemini Director
Our engineering team made great progress recently toward
bringing Gemini-North back on-line after the large earthquake experienced
about 2 weeks ago in Hawaii. As of last Friday the only subsystem
preventing us from going back onto the sky for night-time engineering
tests was the secondary mirror's tip/tilt system, which would not
initialize properly. Recent tests have demonstrated that the problem
is very likely due to a bad position sensor and/or associated cabling.
Unfortunately these components are deeply embedded within the tip/tilt
system and we must remove the secondary mirror from the telescope and
perform the necessary repair work in the summit lab. Given the time
needed to remove the secondary from the telescope, receive new parts,
install them, then mount the secondary mirror assembly back on the
telescope, we are now planning on having Gemini North shutdown for at
least 2 more weeks.
When the secondary mirror is repaired and night-time tests are
possible, we still must make a variety of telescope performance tests
before going back into science operations.
I would like to once again thank all those who have worked
incredibly hard to bring Gemini-North back on-line. We've come a long way
since the earthquake, but clearly have more work ahead.
The following items were updated at 5h00 pm HST on Tuesday, October 24
- Today's work focused on the SCS (Secondary Control System). It appeared the problem was with the reaction mass control on M2. One of the actuator drive cards was replaced and all seemed well. However, once everything was put together the system did not work.
- The laser enclosure cooling problem was resolved today.
- Top shutter was fully opened with no problems as were both sets of wind gates.
- The maximum slew rate of the enclosure has been halved until the pintle bolts on the affected bogies are replaced.
- Tomorrow work will focus on the SCS.
The following items were updated at 5h00 pm HST on Monday, October 23
- Work related to several tasks of verification and adjustment continued on Saturday and today, Monday. The engineering and technical team had a well deserved day off on Sunday.
- Almost all systems appear ready to go back on the sky for engineering tests. However, the Secondary Control System (SCS) is challenging us. The SCS involves a set of very sophisticated servo systems capable of moving the secondary mirror side-ways (X & Y motion), along the optical axis in focus (Z), run fast tip-tilt, and chop at ~3 Hz. Currently we are experimenting a problem with the Z axis motion where odd behavior is observed. No component appears broken, but this motion does not work correctly at the moment. We need to solve this problem to make further progress.
- In parallel, some other tasks related to datuming the mount control system (MCS) have continued. The laser was also powered up briefly.
The following items were updated at 5h00 pm HST on Friday, October 20
After intensive work of inspections, repairs, adjustments and calibrations, we are close to going back on the sky for engineering test. This will allow a full verification of all systems in a dynamic state and demonstrate how close we are to normal operation. Pending good weather (which is not looking good at the moment), we plan to start nighttime engineering on the sky over the next few nights.
- We can now move the telescope fully and at all operational speeds both in azimuth and elevation; more exercise of these movements will done tomorrow;
- While the telescope was pointing at low elevation, we conducted extensive visual inspection of the secondary mirror and of its associated components. M2 looks fine. The mirror was manually unbound from its limit position where its got stuck during the earthquake;
- We are currently trying to activate all SCS system to control M2. This task will continue tomorrow. A fully functional M2 is essential to our return to night time sky observing;
- We also conducted full visual inspection of the laser beam transfer optics and of the Laser Launch Telescope behind the secondary mirror unit. They were found in good shape, with no apparent damage nor gross misalignment;
- The vanes supporting the secondary mirror and system were inspected;
- The covers of all the dome light fixtures were verified to avoid loose units falling;
- Further work was conducted on the telescope enclosure. Tension on the chains of the upper and lower shutter was adjusted;
Our engineering and technical staff have done a stupendous job. They will work tomorrow Saturday, but they will take a day off on Sunday.
The following items were updated at 5h00 pm HST on Thursday, October 19:
- today was a very productive day;
- repair on the Hydrostatic Bearing Control system on the azimuth went very well. The radial bearing was adjusted and the bearing pressure looks good;
- the telescope was moved in azimuth at max speed between +/- 270 degrees;
- the elevation axis is on oil and tomorrow the axial bearings will be adjusted;
- full inspection of M1, mirror covers and M1 baffle. All the dust that was on the mirror cover ended up on M1!
- inspection of the cass rotator shows everting is fine;
- top and bottom shutter work fine;
- still working on the laser enclosure and the center section of the catwalk;
- PWFS1 is being realigned;
- tomorrow the telescope will be moved to the horizon and a full inspection of M2 and the secondary control system will be performed;
The following items were updated at 5h00 pm HST on Wednesday, October 18:
* The azimuth cable wrap was repaired, adjusted and tested.
* The telescope was rotated in azimuth (+/ 90 degrees) using the handle paddle. The tape encoders tracked the telescope movement well. The alignment of the tape encoders will be checked tomorrow.
* The telescope was positioned so that M2 could be inspected from the 'cherry picker'. The visual inspection of M2 showed no visible signs of damage. A great relief to everyone.
* The state of the secondary tip-tilt mechanism is still unknown
* Some of the pintle bolts on six bogies are broken and will be replaced. This is not stopping us from rotating the enclosure.
* GMOS has been checked thoroughly and appears to be operating nominally. The GMOS optics have not shifted relative to the mask plane or the detector.
The following items were updated at 5h00 pm HST on Tuesday October 17:
* Today's main work was centered on the hydrostatic bearing system
(HBS). Oil was injected in the system and the telescope was re-
centered on the azimuthal track. Alignment of the encoders needs to be
checked and a first head will be prepared. An oil spacing of 75
microns implies a good microfilm of oil between the supporting pads of
the telescope and the azimuthal track. The telescope was rotated by
* Inspection of the dome enclosure reveal no structural problem.
However the top and bottom shutters appear to be out of alignment.
* The acquisition and guiding unit (A&G) was tested and mechanism were
moved and found ok.
* The Calibration unit was also checked and everything appears to be in
a fine state.
* Some functionalities of GMOS were tested without problem.
* Chiller #1 and #2 were repaired and are now fully functional. The
enclosure is now cooled as well as all electronic cabinets -- their
doors have been closed.
* The laser bench optics were checked visually: no loose nor broken
parts found. (The laser has not been powered nor tested yet).
* Tomorrow the work will concentrate on the elevation axis.
* It is still impossible to predict when we will return to normal
operation. We have not done even visual inspection to a crucial part of
our system: the secondary mirror and its related systems. We need to be
able to move the telescope to conduct this inspection.
The following items were updated at 5:00 PM (HST) on Monday, October 16:
* Inspection of the Gemini North Mauna Kea building and enclosure took
place today. No major problems were found. Elevator is safe and
has not been affected by the earthquake.
* Glycol and oil pipes have not been damaged and level 5 (telescope
floor) shows no damage.
* Cooling systems appears fine, but the two main chillers are not
operational, each for a different reason
* All cryo-coolers are functioning and instruments are all close to
* The telescope mount has shifted with respect to the azimuthal track.
Bearing system appears fine and oil will be injected tomorrow to
energize the mount. Elevation axis seems to be in a nominal state.
* Primary mirror control system is in good shape. Mirror cover has not be
activated, nor opened. But indirect inspection indicates that the
primary mirror is in one piece; there appears to be some debris on the
surface of the mirror.
* All electronic cabinets are fine and functioning.
* Instruments: temperatures and vacuum are normal. No mechanism nor
component have been activated yet. Optics of GMOS and ALTAIR are fine.
* Inspection will continue tomorrow, with main effort on the mount in
order to enable motion of the telescope. This will allow us to inspect
the secondary mirror and its associated system. This is a crucial step
in our full inspection.
It is impossible to predict when we will return to normal operation.
We estimate that continuing inspection, test, checks and science
verifications will take several days, even if no major stumbling block
Another update will follow when more tests are done on Tuesday, October 17th.
The following items were updated at 9:30 AM on Monday, October 16, 2006:
* Generator power and the UPS systems were activated shortly after the earthquake to restore the helium compressors and instrument cooling systems. This quick action kept the instruments cold and we do not anticipate a lengthy period of time to cryo-cycle them.
* Hawaiian Electric Company (HELCO) restored stable power to the summit at ~2:30 PM HST (Sunday, October 15). We have switched over to HELCO power as part of the process of bringing all systems back on-line.
* Visual inspections of the Mauna Kea control room, telescope structure, and dome indicate no major damage to Gemini North. That said, the observatory clearly shook hard during this event and there is evidence that the telescope moved on the hydrostatic bearing surfaces (w/out oil of course). There is also evidence that the earthquake restraint system may have engaged as well, this being the ultimate fallback to prevent the telescope from tipping over on the main support pier.
* The telescope was closed on Sunday night (October 15). Extensive tests need to be performed to determine if the telescope was damaged and this will require considerable time to complete. We are doing this using a full crew that went up on Monday morning. We need to methodically evaluate the condition of all enclosure, building, telescope and subsystems to make sure we (1) understand the true state of the telescope and (2) don't make matters worse by exercising them prematurely.
* It is likely that it will take several days until we are ready to resume regular science operations at Gemini North assuming that no serious damage is found in our assessments over the next 24-48 hours.
News Archive Filter
The GEMMA Podcast
A podcast about Gemini Observatory and its role in the Era of Multi-Messenger Astronomy. Featuring news related to multi-messenger astronomy (MMA), time-domain astronomy (TDA), our visiting instrument program, and more through interviews with astronomers, engineers, and staff both here at Gemini (North and South) and abroad.